OpenStudy (anonymous):

A cyclist rides his bike at a speed of 27 miles per hour. What is this speed in miles per minute? How many miles will the cyclist travel in 15 minutes? Do not round your answers.

OpenStudy (anonymous):

Anyone know the answer?

OpenStudy (anonymous):

*answers haha

OpenStudy (wolf1728):

27 miles per hour 60 min per hour so 27mph = 27/60 miles per minute = .45 miles per minute

OpenStudy (the_fizicx99):

Hmm, 27 per hour so each minute would be 27/60 = .45 for 15 minutes, 0.45*15 = 6.75

OpenStudy (wolf1728):

seems you've got your answers now :-)

OpenStudy (anonymous):

thanks so much guys! :)

OpenStudy (wolf1728):

u r welcome